Consider a town in which only two residents, Gilberto and Juanita, own wells that produce water safe for drinking. Gilberto and Juanita can pump and sell as much water as they want at no cost. For them, total revenue equals profit. The following table shows the town's demand schedule for water.

Price

Quantity Demanded

Total Revenue

(Dollars per gallon)

(Gallons of water)

(Dollars)

3.60 0 0
3.30 35 $115.50
3.00 70 $210.00
2.70 105 $283.50
2.40 140 $336.00
2.10 175 $367.50
1.80 210 $378.00
1.50 245 $367.50
1.20 280 $336.00
0.90 315 $283.50
0.60 350 $210.00
0.30 385 $115.50
0 420 0

Suppose Gilberto and Juanita form a cartel and behave as a monopolist. The profit-maximizing price is ___ per gallon, and the total output is ___ gallons. As part of their cartel agreement, Gilberto and Juanita agree to split production equally. Therefore, Gilberto's profit is ___, and Juanita's profit is ___.

Suppose that Gilberto and Juanita have been successfully operating as a cartel. They each charge the monopoly price and sell half of the monopoly quantity. Then one night before going to sleep, Gilberto says to himself, "Juanita and I aren't the best of friends anyway. If I increase my production to 35 gallons more than the cartel amount, I can increase my profit even though her profit goes down. I will do that starting tomorrow."

After Gilberto implements his new plan, the price of water ___ to ___ per gallon. Given Juanita and Gilberto's production levels, Gilberto's profit becomes ___ and Juanita's profit becomes ___.

Because Gilberto has deviated from the cartel agreement and increased his output of water to 35 gallons more than the cartel amount, Juanita decides that she will also increase her production to 35 gallons more than the cartel amount.

After Juanita increases her production, Gilberto's profit becomes ___, Juanita's profit becomes ___, and total profit (the sum of the profits of Gilberto and Juanita) is now ___.

True or False: Based on the fact that both Gilberto and Juanita increased production from the initial cartel quantity, you know that the output effect was larger than the price effect at that quantity.

Note that Gilberto and Juanita started by behaving cooperatively. However, once Gilberto decided to cheat, Juanita decided to cheat as well. In other words, Juanita's output decisions are based on Gilberto's actions.

This behavior is an example of ___.
Answer #1

Blanks-

1) 1.80

2) 210

3) 378/2 = 189

4) 378/2 = 189

5) decreases

6) 1.50

7) 140*1.5 = 210

8) 105*1.5 = 157.5

9) 140*1.2 = 168

10) 140*1.2 = 168

11) 168+168 = 336

The statement is True

This is an example of a tit-for-tat strategy